1.Suppose security C pays $800 if the economy is weak and $200 if the economy is strong. Risk free rate is 5% and the market index security (MIS) pay $800 and $1400 is the economy is weak and strong, correspondingly. MIS is priced at $950

a) what is no-arbitrage price of security C?

b) what is the expected return of security C?

c) what is the risk premium of security C?
Answer #1

Answer:

(a).

Let p be the probability of weak econcomy and 1-p be the probability of strong economy

800*p+1400*(1-p)=950*1.05

=>p=0.671

Price of Security C=(800*0.671+200*(1-0.671))/1.05=573.81